I have 3 side measures given to me, but how do I use trig to find the angle measures?
Answer by cleomenius(959)

If this is a right triangle, you can use the tangent function, tan angle = opp/adj.
Once you have one angle, you can use the law of sines to find another angle.
sin A/ side A = Sin B/ side b.
Once you have the measure of two angles, you can find the remaining angle by adding the two and subtracting from 180.
=============================================================================
If it is not a right triangle, you can use the law of Cosines to find angle C.
You should find the largest angle first, because there can only be one angle in a triangle that is obtuse. The reason for finding it first is the next step, the sine rule, can only give us acute angles ( less than 90 degrees), so we avoid a possible wrong answer by first eliminating the only possibility of an obtuse angle.
c^2 = b^2 + a^2 = 2bacosC
cosC = a^2 + b^2 - 2abcosC

cosC = a^2 + b^2 -c^2/2ab
From here, you can also obtain a second angle using the law of sines and third angle by adding the two and subtracting from 180.
Cleomenius.